OpenStudy (cheesenips111111):

What is the value of z in the equation 2z + 6 = −4? 5 1 −1 −5

OpenStudy (cheesenips111111):

@TheSmartOne

TheSmartOne (thesmartone):

Sorry, got distracted on another tab :P So what did you get as the answer?

OpenStudy (cheesenips111111):

5?

OpenStudy (cheesenips111111):

@TheSmartOne

TheSmartOne (thesmartone):

What did you get for -4 - 6 = ?

OpenStudy (cheesenips111111):

no i think i got the answer

OpenStudy (cheesenips111111):

5

OpenStudy (cheesenips111111):

or a

TheSmartOne (thesmartone):

Yeah, that's wrong :P -4-6 = -10 and then if you divide by 2, you get a negative answer, not a positive one

OpenStudy (cheesenips111111):

oh so its -5?

TheSmartOne (thesmartone):

Yes

OpenStudy (cheesenips111111):

ok i have just a couple more

OpenStudy (cheesenips111111):

if that s ok?

OpenStudy (cheesenips111111):

The incorrect work of a student to solve an equation 2(y + 8) = 4y is shown below: Step 1: 2(y + 8) = 4y Step 2: 2y + 10 = 4y Step 3: 2y = 10 Step 4: y = 5 Which of the following explains how to correct Step 2 and shows the correct value of y? The equation should be y + 8 = 4y after division by 2; y = 4 The equation should be y + 8 = 4y after division by 2; y = 6 2 should be distributed as 2y + 16; y = 4 2 should be distributed as 2y + 16; y = 8

OpenStudy (cheesenips111111):

i got d

TheSmartOne (thesmartone):

correct :)

OpenStudy (cheesenips111111):

ok

OpenStudy (cheesenips111111):

Which set of steps shows the solution to the equation 3y = −9? y = −9 − 3; y = 6 y = −9 ÷ 3; y = −3 y = −9 ÷ (−3); y = 3 y = −9 − (−3); y = −6

OpenStudy (cheesenips111111):

b

TheSmartOne (thesmartone):

correct :)

OpenStudy (cheesenips111111):

ok

OpenStudy (cheesenips111111):

A system of equations is shown below: n = 3m + 7 n − 2m = 1 What is the solution, in the form (m, n), to the system of equations? (3, 7) (1, 8) (−3, −2) (−6, −11)

OpenStudy (cheesenips111111):

b

OpenStudy (cheesenips111111):

c

OpenStudy (cheesenips111111):

i mean

OpenStudy (cheesenips111111):

i mean C

TheSmartOne (thesmartone):

That's wrong. How do you want to solve it? Elimination or substitution?

OpenStudy (cheesenips111111):

ok

OpenStudy (cheesenips111111):

is it b?

TheSmartOne (thesmartone):

Nope. So which one will it be? Elimination or substitution?

OpenStudy (cheesenips111111):

elimination?

TheSmartOne (thesmartone):

n = 3m + 7 n − 2m = 1 let's add 2m on both sides to the second equation n = 3m + 7 n = 2m + 1 what do you get when you subtract the two equations?

OpenStudy (cheesenips111111):

A? right

OpenStudy (cheesenips111111):

??

OpenStudy (cheesenips111111):

you there i have to check this by 9

OpenStudy (cheesenips111111):

i hate bieng annoying

TheSmartOne (thesmartone):

You're just guessing random answers now.

OpenStudy (cheesenips111111):

idk what it is but ive got to get this done before tonit because my dad wont sign me up for lacrosse if i dont he wont

TheSmartOne (thesmartone):

That's alright, but the question won't take more than 2 minutes if you cooperate with me. n = 3m + 7 n = 2m + 1 Subtract the two equations. If we do it right now, we'll be getting rid of n because n - n = 0. And we'll then be left with m and find the value of m. so, let's subtract n - n = 3m + 7 - (2m + 1) simplify that :)

OpenStudy (cheesenips111111):

its D (-6,-11)

OpenStudy (cheesenips111111):

right i simplyfid i swear

TheSmartOne (thesmartone):

Yup! Good job :)

OpenStudy (cheesenips111111):

ok

OpenStudy (cheesenips111111):

The equation 1.6x − 2.2 = 1.4x can be transformed to form which of the following expressions? 3.0x = 2.2 0.2x = −2.2 3.0x = −2.2 0.2x = 2.2

OpenStudy (cheesenips111111):

i got D

TheSmartOne (thesmartone):

correct

OpenStudy (cheesenips111111):

ok

OpenStudy (cheesenips111111):

A system of two equations is shown below: Equation C: a = 3b + 6 Equation D: a = 7b − 1 What value of a can be substituted into equation D to solve the system of equations? 3b 7b 3b + 6 7b − 1

OpenStudy (cheesenips111111):

C?

TheSmartOne (thesmartone):

Correct! :)

OpenStudy (cheesenips111111):

ok

OpenStudy (cheesenips111111):

Which statement is true for the equation 3x − 3x − 2 = −2? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

OpenStudy (cheesenips111111):

D?

OpenStudy (cheesenips111111):

?

TheSmartOne (thesmartone):

Correct :)

OpenStudy (cheesenips111111):

yay

TheSmartOne (thesmartone):

:D

OpenStudy (cheesenips111111):

The system of equations shown below is graphed on a coordinate grid: 4y + x = 4 2y − x = 8 Which statement is true about the coordinates of the point that is the solution to the system of equations? It is (2, 5) and lies on both lines. It is (−4, 2) and lies on both lines. It is (2, 5) and does not lie on either line. It is (−4, 2) and does not lie on either line.

OpenStudy (cheesenips111111):

A?

TheSmartOne (thesmartone):

When you do elimination, you're solving for the variable y not x so you get y = 2, not x = 2 so try again :)

OpenStudy (cheesenips111111):

B?

OpenStudy (cheesenips111111):

@TheSmartOne

TheSmartOne (thesmartone):

correct :)

OpenStudy (cheesenips111111):

ok

OpenStudy (cheesenips111111):

Which statement is true for the equation 5n − 4 = 5n − 3? It has infinitely many solutions. It has two solutions. It has one solution. It has no solution.

OpenStudy (cheesenips111111):

D?

TheSmartOne (thesmartone):

correct :)

OpenStudy (cheesenips111111):

Which statement is true about the solutions for the equation 4y + 6 = −2? It has infinitely many solutions. It has two solutions. It has one solution. It has no solution.

OpenStudy (cheesenips111111):

c?

TheSmartOne (thesmartone):

bingo (:

OpenStudy (cheesenips111111):

yay

OpenStudy (cheesenips111111):

Which graph best represents the solution to the system of equations shown below? y = -4x - 19 y = 2x − 1 A coordinate grid is shown from negative 10 to positive 10 on the x axis and also on the y axis. Two lines are shown intersecting on ordered pair 7, 3. A coordinate grid is shown from negative 10 to positive 10 on the x axis and also on the y axis. Two lines are shown intersecting on ordered pair 3, 7. A coordinate grid is shown from negative 10 to positive 10 on the x axis and also on the y axis. Two lines are shown intersecting on ordered pair negative 3, negative 7. A coordinate grid is shown from negative 10 to positive 10 on the x axis and also on the y axis. Two lines are shown intersecting on ordered pair negative 7, negative 3.

OpenStudy (cheesenips111111):

D?

TheSmartOne (thesmartone):

Nope, you have the x and y coordinates switched

OpenStudy (cheesenips111111):

A?

OpenStudy (cheesenips111111):

@TheSmartOne

TheSmartOne (thesmartone):

nope y = -4x - 19 y = 2x − 1 subtract both equations y - y = -4x - 19 - (2x - 1) simplify it :)

OpenStudy (cheesenips111111):

BBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBbb

OpenStudy (cheesenips111111):

B

OpenStudy (cheesenips111111):

(;

OpenStudy (cheesenips111111):

right????

TheSmartOne (thesmartone):

nope, and by process of elimination, you now know what the answer is

OpenStudy (cheesenips111111):

A pencil at a stationery store costs $1, and a pen costs $1.50. Stefan spent $7 at the store. He bought a total of 6 items. Which system of equations can be used to find the number of pencils (x) and pens (y) he bought? x + 1.5y = 7 x + y = 6 1.5x + y = 7 x = 6y x + 2y = 7 6x = 1.5y 2x + y = 7 6x = 1.5y

OpenStudy (cheesenips111111):

last one and A

TheSmartOne (thesmartone):

correct :)

OpenStudy (cheesenips111111):

yay

OpenStudy (cheesenips111111):

thank you soooo much you helped sooo much

TheSmartOne (thesmartone):

Anytime! :)